OpenStudy (anonymous):

Please Help! I always give medals. A Hummer four-door costs the dealer $57,915. The dealer has a 12.3% markup on selling price. What is the selling price of the vehicle? Round to the nearest dollar. Note: Markup is on selling price, not cost.

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

add 12.3% of the cost, back to the cost ...

OpenStudy (amistre64):

percents are just a division by 100 for a given number n, 1% is equal to n/100 to find say, 12%, we want to know the value of 12 times n/100

OpenStudy (amistre64):

they give us: 57,915 1% = 57,915/100 12.3% = 12.3 * 57,915/100 just add that amount back into 57,915

OpenStudy (anonymous):

is it $7,123.55 ?

OpenStudy (anonymous):

yes maybe

OpenStudy (anonymous):

So that would be $65,038.55 rounded $65,039 ? is that correct?